Hydrogen gas is generated by the decomposition of water to fill a 1.1 kL weather balloon at 20°C and 100.0 kPa. What is the mass of hydrogen required?
1
Expert's answer
2020-06-08T15:36:03-0400

P = 100.0 kPa

V = 1.1 kL = 1100 L

T = 20°C = 293 K

R = 8.314 kPa L mo-1 K-1


Solution:

The decomposition reaction of water:

2H2O → 2H2 + O2


Ideal Gas Law can be used.

Ideal Gas Law can be expressed as: PV=nRT

To find the moles of H2, solve the equation for n:

n = PV/RT

Moles of H2 = n(H2) = (100.0 kPa × 1100 L) / (8.314 kPa L mol-1 K-1 × 293 K) = 45.156 mol


Moles of H2 = Mass of H2 / Molar mass of H2

The molar mass of H2 is 2.016 g/mol.

Finally, the mass of H2 is:

m(H2) = n(H2) × M(H2) = (45.156 mol) × (2.016 g/mol) = 91.034 g = 91 g

The mass of H2 is 91.0 g


Answer: 91 g is the mass of hydrogen required.